OpenStudy (anonymous):

A rod 7.5 ft long is to be cut into 5 pieces. How many feet long will each piece be?

OpenStudy (anonymous):

Nancy has a sewing project which calls for panels which use 1 1/4 yd of fabric each. She has 4 1/2 yd of fabric. How many panels can she cut from the fabric? A. 2 B. 3 C. 4 D. 5

OpenStudy (anonymous):

B. 3 \[\frac{4+\frac{1}{2}}{1+\frac{1}{4}}=3\frac{6}{10} \]